# Brain3D: Generating 3D Objects from fMRI

### v4

@@ 4.4 Brain region functionalities
Specifically, we observe that the two hemispheres make different holistic contributions to object perception, due to exhibiting distinct performances for specific objects (Fig. 5b). The left hemisphere tends to depict finer details and intricate structures in object visualization, while the right hemisphere better captures the overall shape and silhouette. These results align with previous discoveries that the left hemisphere is involved in detail-oriented tasks while the right hemisphere in more holistic tasks [67]. However, our results challenge the notion of strict functional division. We found notable overlaps and interactions between the two hemispheres. Remarkably, numerous objects rendered by the right hemisphere, as illustrated in Fig. 5a, also exhibit intricate textures—typically a characteristic attributed to the left hemisphere. Conversely, the left hemisphere demonstrated capability in detecting overarching shapes, which is usually a strength of the right hemisphere.
Despite these different functional tendencies, our quantitative analysis, using metrics like 3D Inception, 3D CLIP, 3D EfficientNet, and 3D Swav in Fig. 5c, shows that both hemispheres contribute similarly to 3D object perception. Their collaboration significantly contributes to the quality of 3D objects (Fig. 5c). This result echos the biological discovery on the interaction between the left and right brain [68, 67].
@@
Previous studies have demonstrated that visual cortex [69] is predominantly involved in processing color and texture information, whereas the medial temporal lobe (MTL) region [70, 71] is more focused on semantic information. In comparison, our research shows that the visual cortex exhibits higher visual decoding ability than MTL region (Fig. 6b). The significance level refers to the probability of error when rejecting the null hypothesis that visual cortex shares the similar performance as MTL region. Specifically, the visual cortex captures local fine-grained visual information such as the “windows” of an airplane and the “motorcycle rider”, while the MTL region captures holistic shapes like “airplane skeleton” and “person” (Fig. 6d). This visualization highlights higher-level 3D-shape-related concepts while foregoing pixel-level details. Our extensive analysis illustrated in Fig. 6f shows a high level of semantic consistency along with considerable texture and geometric consistency. This discrepancy may be attributed to the inherent nature of human visual processing, where the brain may not emphasize all texture details, particularly in background or less prominent areas. In regional analysis, we observe that hemispheric symmetry is evident as the left and right hemispheres exhibit similar scores across texture, semantic, and geometric metrics (Fig. 6g), indicating balanced processing. Furthermore, significant visual cortex dominance is apparent in perceptual processes (Fig. 6h), demonstrating a superior capability to integrate geometry, semantics, and textures compared to the MTL, which, while showing reasonable semantic consistency, lacks consistency in geometry and texture. This finding offers a novel perspective on how the MTL may contribute to the formation and recall of long-term memories, highlighting its role in abstracting complex visual information at the semantic level.
@@
In line with earlier researches [7], V1 region exhibits predominant ability in visual representation (Fig. 6c) among various visual regions. To further break down, V1 region plays a pivotal role in handling elementary visual information, including edges, features, and color. The V1 region’s essential function in managing elementary visual information is underscored by its decoding results (Fig. 6i), which reveal more substantial semantic, geometric, and texture consistency over other regions. V2 and V3 regions play a role in synthesizing various features into complex visual forms, while V4 region focuses on basic color process at coarse level (Fig. 6e). However, our research brings to light a nuanced understanding of the interdependence within the visual cortex. Specifically, we demonstrate the reliance of the V2, V3, and V4 regions on the V1 region for constructing 3D vision. This is evident from the disrupted 3D geometries observed in Fig. 6e. Moreover, scenes decoded from the V2, V3, and V4 regions exhibit only moderate consistency in texture, semantics, and geometry (Fig.6g). This cross-region reliance likely stems from the fact that these regions process information initially received by V1 [3], indicating a foundational dependence on V1 for effective visual processing. These findings emphasize the collaborative and interactive roles of these brain regions in forming a comprehensive visual representation. This result is thus biologically consistent to great extent.
@@ 4.5 Clinical evaluation
To simulate the disorders of varying brain regions, we adopt the general fMRI simulation techniques [72] that introduce Gaussian noises. We randomly select samples from the NSD test set for disorder simulation. For each sample, we apply independent Gaussian noises with a variance of 3.0 over five brain regions: V1, V2, V3, V4, and MTL, respectively.
@@
Visual region clinical diagnosis Motivated by the high biological consistency in terms of regional functions, we examine the usefulness of Brain3D in diagnosing disordered brain regions by designing a series of simulation based experiments (Fig. 7a). To compare performance of Brain3D and 2D models, we designed three branches of experiments: 2D model MindEye [18], Brain3D front view and Brain3D back view. In each branch, the participants were tasked with determining the impacted brain region of the patient from V1, V2, V3, V4, and MTL by observing 3D visuals generated from their fMRI signals. Their average correctness of judgement is represented as evaluation of clinical diagnosis. Notably, almost perfect diagnosis is achieved for the disorders with the V1 and MTL regions (Fig. 7b). However, distinguishing the disorders among V2, V3, and V4 regions presents more challenges, as reflected by the incidence of incorrect diagnoses in these areas. This difficulty is attributed to the significant overlap and interaction in their processing roles, as they all engage in visual information processing following the V1 region. Nonetheless, when a disorder was localized to a single region among V2, V3, or V4, participants generally succeeded in identifying that the issue stemmed from one of these three regions, despite the nuanced differences in their functionalities. Diagnosis among all regions in Brain3D outperforms previous 2D methods (Fig. 7c). We provide visual examples for region disorder diagnosis (Fig. 7d). V1 disorder tends to cause significant distortion to the output, consistent with the earlier finding about its critical role in processing basic visual information. Disorder in V2 region leads to changes in the detailed shape and motion of objects, e.g., orientation of the airplane and shape of the motorcycle. V3 region primarily affects the details and movements, e.g., the change from motorcycle to rider. Disorder in V4 region might cause color shift and variation. Instead, MTL’s disorder could be concerned with high-level semantics, e.g., increased ambiguity in object appearance. See Sec. 4.5 for more details.
@@
Diagnosis on complex simulations To explore diagnosis through more complex neurological conditions, we further extend our experiments with Rician [73, 74] (with a variance of 3.0), Drift [75] (with a drift rate of 0.1), and Physiological [76, 77] (with a fluctuation intensity of 0.5) simulations across all five brain regions (V1, V2, V3, V4, and MTL). Both Drift and Physiological simulations are further combined with Gaussian noise to mimic more complex neurological conditions [75, 76, 77]. Figure 9(a) shows how different simulated disorders impact brain regions. For V1, each simulation type creates distinct distortions: Rician makes results vague, Gaussian distorts shapes, Drift breaks up textures, and Physiological adds irrelevant features. Despite these variations, all V1 simulations severely disrupt basic visual information. In V2, all simulations lead to changes in detailed shapes. For V3, they slightly alter fine details. For V4, disorders primarily change detailed colors, like making the airplane whiter and the clock tower bluer. Importantly, even with these changes, the overall outlines of objects like airplanes and clock towers remain largely visible across V2, V3, and V4, highlighting these regions’ supportive roles. Disorders in the MTL region, across all simulations, increase the ambiguity of object appearance, showing its link to high-level meaning. Figure 9(b) confirms that our diagnostic method provides reliable conclusions across all simulations.
@@ 4.6 Ablation analysis
Ablation on guidance Our perceptual pipeline employs Score Distillation Sampling (SDS) [47] with the inclusion of both low-level and high-level guidance. The integration of both low-level and high-level guidance achieves the best performance across all categories (Fig. 10a). As shown in Fig. 10b, when solely relying on high-level guidance, the representations of the train, airplane, and clock tower appear significantly blurred, displaying only coarse shapes. These coarse shapes stem from the geometric and semantic information embedded within the high-level representation. Low-level guidance primarily captures the color and texture attributes of the objects. Consequently, the outcomes with low-level guidance (see the first column), featuring the front view, align well with the provided image, while other views fall short in terms of shape and semantic fidelity. The combination of both high-level and low-level guidance results in significantly improved representations of the airplane, train, and clock tower, where both the semantic and textural structures closely resemble those of the testing image. Ablation on the Semantic Stage The Semantic Stage within our framework is designed to extract finer details from fMRI embeddings, enhancing its contribution to 3D generation. After the Semantic Stage the detailed textures of the building in the first row and the outline of the steamship in the second row in Fig. 11c exhibit much greater clarity. These details from fMRI are challenging to represent in a coarse NeRF but can be effectively distilled within a 3D mesh.
@@
Ablation on visual stimuli To explore the purely imaginative scenario of generating objects without relying on reference images, we leveraged the TMNRED dataset [78]. This dataset is particularly well-suited for this purpose as participants’ brain signals were collected exclusively under text description stimuli, with no accompanying visual input. Our Brain3D model is designed to process generic brain signals (e.g., EEG), regardless of the format of the visual stimuli used during data collection. As demonstrated in Figure 13(a), our model successfully generates visuals that exhibit a commendable alignment with the textual prompts, even when conditioned solely on brain signals derived from text-based imagination and in the complete absence of image stimuli. This capability underscores our model’s unique ability to directly decode and reconstruct mental imagery from brain activity, highlighting its potential to tap into the human imagination guided by text, past experience, and knowledge. Furthermore, we investigated the effect of incorporating a synthetic visual prior. When our model was additionally equipped with a paired training image generated by Stable Diffusion [48] (conditioned on the original text), Figure 13(b) shows that finer object details can be recovered. This suggests that our model can effectively integrate such a generative visual prior to enhance the translation from fMRI signals to more detailed and refined visual representations.
@@ 6 Data Availability Statement
The datasets generated and/or analysed for modeling during the current study are available in: Natural Scenes Dataset (NSD) [19] (https://natural-scenes-dataset.s3.amazonaws.com/index.html). The dataset generated and analyzed in brain region analysis and diagnosis will be made available upon request after the peer-review process. Interested researchers can request access by contacting the corresponding author via the journal’s communication channels.

### v5

@@ 4.4 Brain region functionalities
Specifically, we observe that the two hemispheres make different holistic contributions to object perception, due to exhibiting distinct performances for specific objects (Fig. 5b). The left hemisphere tends to depict finer details and intricate structures in object visualization, while the right hemisphere better captures the overall shape and silhouette. These results align with previous discoveries that the left hemisphere is involved in detail-oriented tasks while the right hemisphere in more holistic tasks Iaccino (2014). However, our results challenge the notion of strict functional division. We found notable overlaps and interactions between the two hemispheres. Remarkably, numerous objects rendered by the right hemisphere, as illustrated in Fig. 5a, also exhibit intricate textures—typically a characteristic attributed to the left hemisphere. Conversely, the left hemisphere demonstrated capability in detecting overarching shapes, which is usually a strength of the right hemisphere.
Despite these different functional tendencies, our quantitative analysis, using metrics like 3D Inception, 3D CLIP, 3D EfficientNet, and 3D Swav in Fig. 5c, shows that both hemispheres contribute similarly to 3D object perception. Their collaboration significantly contributes to the quality of 3D objects (Fig. 5c). This result echos the biological discovery on the interaction between the left and right brain Corballis (2014); Iaccino (2014).
@@
Previous studies have demonstrated that visual cortex Milner and Goodale (2006) is predominantly involved in processing color and texture information, whereas the medial temporal lobe (MTL) region Squire et al. (2004); Eichenbaum et al. (2007) is more focused on semantic information. In comparison, our research shows that the visual cortex exhibits higher visual decoding ability than MTL region (Fig. 6b). The significance level refers to the probability of error when rejecting the null hypothesis that visual cortex shares the similar performance as MTL region. Specifically, the visual cortex captures local fine-grained visual information such as the “windows” of an airplane and the “motorcycle rider”, while the MTL region captures holistic shapes like “airplane skeleton” and “person” (Fig. 6d). This visualization highlights higher-level 3D-shape-related concepts while foregoing pixel-level details. Our extensive analysis illustrated in Fig. 6f shows a high level of semantic consistency along with considerable texture and geometric consistency. This discrepancy may be attributed to the inherent nature of human visual processing, where the brain may not emphasize all texture details, particularly in background or less prominent areas. In regional analysis, we observe that hemispheric symmetry is evident as the left and right hemispheres exhibit similar scores across texture, semantic, and geometric metrics (Fig. 6g), indicating balanced processing. Furthermore, significant visual cortex dominance is apparent in perceptual processes (Fig. 6h), demonstrating a superior capability to integrate geometry, semantics, and textures compared to the MTL, which, while showing reasonable semantic consistency, lacks consistency in geometry and texture. This finding offers a novel perspective on how the MTL may contribute to the formation and recall of long-term memories, highlighting its role in abstracting complex visual information at the semantic level.
@@
In line with earlier researches Tong (2003), V1 region exhibits predominant ability in visual representation (Fig. 6c) among various visual regions. To further break down, V1 region plays a pivotal role in handling elementary visual information, including edges, features, and color. The V1 region’s essential function in managing elementary visual information is underscored by its decoding results (Fig. 6i), which reveal more substantial semantic, geometric, and texture consistency over other regions. V2 and V3 regions play a role in synthesizing various features into complex visual forms, while V4 region focuses on basic color process at coarse level (Fig. 6e). However, our research brings to light a nuanced understanding of the interdependence within the visual cortex. Specifically, we demonstrate the reliance of the V2, V3, and V4 regions on the V1 region for constructing 3D vision. This is evident from the disrupted 3D geometries observed in Fig. 6e. Moreover, scenes decoded from the V2, V3, and V4 regions exhibit only moderate consistency in texture, semantics, and geometry (Fig.6g). This cross-region reliance likely stems from the fact that these regions process information initially received by V1 Hubel and Wiesel (1979), indicating a foundational dependence on V1 for effective visual processing. These findings emphasize the collaborative and interactive roles of these brain regions in forming a comprehensive visual representation. This result is thus biologically consistent to great extent.
@@ 4.5 Clinical evaluation
To simulate the disorders of varying brain regions, we adopt the general fMRI simulation techniques Welvaert and Rosseel (2014) that introduce Gaussian noises. We randomly select samples from the NSD test set for disorder simulation. For each sample, we apply independent Gaussian noises with a variance of 3.0 over five brain regions: V1, V2, V3, V4, and MTL, respectively.
@@
Visual region clinical diagnosis Motivated by the high biological consistency in terms of regional functions, we examine the usefulness of Brain3D in diagnosing disordered brain regions by designing a series of simulation based experiments (Fig. 7a). To compare performance of Brain3D and 2D models, we designed three branches of experiments: 2D model MindEye Scotti et al. (2024), Brain3D front view and Brain3D back view. In each branch, the participants were tasked with determining the impacted brain region of the patient from V1, V2, V3, V4, and MTL by observing 3D visuals generated from their fMRI signals. Their average correctness of judgement is represented as evaluation of clinical diagnosis. Notably, almost perfect diagnosis is achieved for the disorders with the V1 and MTL regions (Fig. 7b). However, distinguishing the disorders among V2, V3, and V4 regions presents more challenges, as reflected by the incidence of incorrect diagnoses in these areas. This difficulty is attributed to the significant overlap and interaction in their processing roles, as they all engage in visual information processing following the V1 region. Nonetheless, when a disorder was localized to a single region among V2, V3, or V4, participants generally succeeded in identifying that the issue stemmed from one of these three regions, despite the nuanced differences in their functionalities. Diagnosis among all regions in Brain3D outperforms previous 2D methods (Fig. 7c). We provide visual examples for region disorder diagnosis (Fig. 7d). V1 disorder tends to cause significant distortion to the output, consistent with the earlier finding about its critical role in processing basic visual information. Disorder in V2 region leads to changes in the detailed shape and motion of objects, e.g., orientation of the airplane and shape of the motorcycle. V3 region primarily affects the details and movements, e.g., the change from motorcycle to rider. Disorder in V4 region might cause color shift and variation. Instead, MTL’s disorder could be concerned with high-level semantics, e.g., increased ambiguity in object appearance. See Sec. 4.5 for more details.
@@
Diagnosis on complex simulations To explore diagnosis through more complex neurological conditions, we further extend our experiments with Rician Wegmann et al. (2017); Gudbjartsson and Patz (1995) (with a variance of 3.0), Drift Gupta et al. (2022) (with a drift rate of 0.1), and Physiological Biswal et al. (1996); Madl et al. (2022) (with a fluctuation intensity of 0.5) simulations across all five brain regions (V1, V2, V3, V4, and MTL). Both Drift and Physiological simulations are further combined with Gaussian noise to mimic more complex neurological conditions Gupta et al. (2022); Biswal et al. (1996); Madl et al. (2022). Figure 9(a) shows how different simulated disorders impact brain regions. For V1, each simulation type creates distinct distortions: Rician makes results vague, Gaussian distorts shapes, Drift breaks up textures, and Physiological adds irrelevant features. Despite these variations, all V1 simulations severely disrupt basic visual information. In V2, all simulations lead to changes in detailed shapes. For V3, they slightly alter fine details. For V4, disorders primarily change detailed colors, like making the airplane whiter and the clock tower bluer. Importantly, even with these changes, the overall outlines of objects like airplanes and clock towers remain largely visible across V2, V3, and V4, highlighting these regions’ supportive roles. Disorders in the MTL region, across all simulations, increase the ambiguity of object appearance, showing its link to high-level meaning. Figure 9(b) confirms that our diagnostic method provides reliable conclusions across all simulations.
@@ 4.6 Ablation analysis
Ablation on guidance Our perceptual pipeline employs Score Distillation Sampling (SDS) Poole et al. (2022) with the inclusion of both low-level and high-level guidance. The integration of both low-level and high-level guidance achieves the best performance across all categories (Fig. 10a). As shown in Fig. 10b, when solely relying on high-level guidance, the representations of the train, airplane, and clock tower appear significantly blurred, displaying only coarse shapes. These coarse shapes stem from the geometric and semantic information embedded within the high-level representation. Low-level guidance primarily captures the color and texture attributes of the objects. Consequently, the outcomes with low-level guidance (see the first column), featuring the front view, align well with the provided image, while other views fall short in terms of shape and semantic fidelity. The combination of both high-level and low-level guidance results in significantly improved representations of the airplane, train, and clock tower, where both the semantic and textural structures closely resemble those of the testing image. Ablation on the Semantic Stage The Semantic Stage within our framework is designed to extract finer details from fMRI embeddings, enhancing its contribution to 3D generation. After the Semantic Stage the detailed textures of the building in the first row and the outline of the steamship in the second row in Fig. 11c exhibit much greater clarity. These details from fMRI are challenging to represent in a coarse NeRF but can be effectively distilled within a 3D mesh.
@@
Ablation on visual stimuli To explore the purely imaginative scenario of generating objects without relying on reference images, we leveraged the TMNRED dataset Bai et al. (2025). This dataset is particularly well-suited for this purpose as participants’ brain signals were collected exclusively under text description stimuli, with no accompanying visual input. Our Brain3D model is designed to process generic brain signals (e.g., EEG), regardless of the format of the visual stimuli used during data collection. As demonstrated in Figure 13(a), our model successfully generates visuals that exhibit a commendable alignment with the textual prompts, even when conditioned solely on brain signals derived from text-based imagination and in the complete absence of image stimuli. This capability underscores our model’s unique ability to directly decode and reconstruct mental imagery from brain activity, highlighting its potential to tap into the human imagination guided by text, past experience, and knowledge. Furthermore, we investigated the effect of incorporating a synthetic visual prior. When our model was additionally equipped with a paired training image generated by Stable Diffusion Rombach et al. (2022) (conditioned on the original text), Figure 13(b) shows that finer object details can be recovered. This suggests that our model can effectively integrate such a generative visual prior to enhance the translation from fMRI signals to more detailed and refined visual representations.
@@ 6 Data Availability Statement
The datasets generated and/or analysed for modeling during the current study are available in: Natural Scenes Dataset (NSD) Allen et al. (2022) (https://natural-scenes-dataset.s3.amazonaws.com/index.html). The dataset generated and analyzed in brain region analysis and diagnosis will be made available upon request after the peer-review process. Interested researchers can request access by contacting the corresponding author via the journal’s communication channels.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).